### v2

# A Classification Theorem on Non-compact Embeddings between Besov Spaces

Abstract.
We analyze the embedding properties between Besov spaces, defined on the total space $\mathbb{R}^{n}$ and on bounded domains. We give a complete classification on whether or not these embedding maps satisfy certain weak compactness characterized by the so-called strictly and finitely strictly singular condition. The result extends the recent findings on Sobolev embeddings by offering a refined description of the quality of non‐compactness in the setting of Besov spaces.
2020 Mathematics Subject Classification:
47B06 (primary) 46E35 and 46B45 (secondary)
1. Introduction
It is well known that Sobolev embeddings on domains (say a bounded Lipschitz domain $\Omega\subset\mathbb{R}^{n}$) into the optimal Lebesgue space and the optimal Lorentz space
$W^{k,p}(\Omega)\hookrightarrow L^{\frac{np}{n-kp}}(\Omega)\quad\text{and}\quad W^{k,p}(\Omega)\hookrightarrow L^{\frac{np}{n-kp},p}(\Omega),$
where $1\leq k<n$ and $p\in[1,n/k)$ are non-compact. Given that $L^{\frac{np}{n-kp}}(\Omega)\subsetneq L^{\frac{np}{n-kp},p}(\Omega)$, one might expect a significant difference in the “quality” of non-compactness for these embeddings. However, standard measures of non-compactness, such as the decay of entropy numbers or Kolmogorov numbers, fail to differentiate between these embeddings, as both are maximally non-compact (i.e., the measure of non-compactness coincides with the norm of the embeddings, see [LMOP21]).
In [LM23], it has been observed that the first Sobolev embedding is finitely strictly singular (i.e., the Bernstein numbers decay to 0), while the second Sobolev embedding is not strictly singular, with all Bernstein numbers equal to the norm of the embedding. This observation suggests that strict singularity and finitely strict singularity could provide a more refined distinction between non-compact embeddings, offering a better description of the “quality” of non-compactness comparable with the measure of non-compactness.
The above observations regarding Sobolev embeddings naturally lead to the question of studying the “quality” of non-compactness for embeddings between other function spaces.
In this paper, we consider the embeddings between Besov spaces
$\mathscr{B}_{p_{0}q_{0}}^{s_{0}}\subset\mathscr{B}_{p_{1}q_{1}}^{s_{1}},\mbox{ both on bounded domains and on }\mathbb{R}^{n}.$
We provide a comprehensive and complete classification of the “quality” of non-compactness for the embeddings between Besov spaces in the context of strict singularity and finitely strict singularity, as demonstrated in the following two theorems which are the main results of this paper.
For embeddings between Besov spaces on bounded domains, we obtain:
Theorem 1.1.
Let $p_{0},p_{1},q_{0},q_{1}\in(0,\infty]$, $s_{0},s_{1}\in\mathbb{R}$. Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz open subset. Consider the embedding $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$.
  1. (i)
    $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\not\subset\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$, i.e. there is no embedding, if and only if:
    • $s_{0}-s_{1}<\max(0,\frac{n}{p_{0}}-\frac{n}{p_{1}})$; or
    • $s_{0}-s_{1}=\max(0,\frac{n}{p_{0}}-\frac{n}{p_{1}})$ and $q_{0}>q_{1}$.
  2. (ii)
    The embedding is compact if and only if $s_{0}-s_{1}>\max(0,\frac{n}{p_{0}}-\frac{n}{p_{1}})$.
  3. (iii)
    The embedding is non-compact but finitely strictly singular if and only if:
    • $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$ and $q_{0}<q_{1}$; or
    • $s_{0}-s_{1}=0$, $p_{1}<p_{0}=\infty$ and $q_{0}<q_{1}$.
  4. (iv)
    The embedding is not finitely strictly singular but strictly singular if and only if:
    • $s_{0}-s_{1}=0$, $p_{1}=p_{0}=\infty$ and $q_{0}<q_{1}$; or
    • $s_{0}-s_{1}=0$, $p_{1}\leq p_{0}<\infty$ and $q_{0}<q_{1}$.
  5. (v)
    The embedding is not strictly singular if and only if $s_{0}-s_{1}=\max(0,\frac{n}{p_{0}}-\frac{n}{p_{1}})$ and $q_{0}=q_{1}$.
Remark.
By applying the standard partition of unity argument, the same results in Theorems 1.1 hold if we replace $\Omega$ by a compact smooth manifold. We refer to [Tri92, Chapter 7] for the definition of spaces and leave the details to readers.
For embeddings between Besov spaces on $\mathbb{R}^{n}$, we obtain:
Theorem 1.2.
Let $p_{0},p_{1},q_{0},q_{1}\in(0,\infty]$, $s_{0},s_{1}\in\mathbb{R}$. Consider the embedding $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
  1. (i)
    $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\not\subset\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$, i.e. there is no embedding, if and only if one of the following occurs:
    • $p_{0}>p_{1}$;
    • $s_{0}-s_{1}<\frac{n}{p_{0}}-\frac{n}{p_{1}}$;
    • $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}\geq 0$ and $q_{0}>q_{1}$.
  2. (ii)
    The embedding is non-compact but finitely strictly singular if and only if one of the following:
    • $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$ and $q_{0}<q_{1}$;
    • $s_{0}-s_{1}>\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$.
  3. (iii)
    The embedding is not strictly singular if and only if one of the following occurs:
    • $s_{0}-s_{1}>0$ and $p_{0}=p_{1}$;
    • $s_{0}-s_{1}=0$, $p_{0}=p_{1}$ and $q_{0}\leq q_{1}$;
    • $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$ and $q_{0}=q_{1}$.
We also include a result of embedding of homogeneous Besov spaces $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$, see Theorem 5.1.
In Theorems 1.1 and 1.2, the case where the map is unbounded and the case where it is compact are both well-known. The critical exponents cases, which form the core of our paper, fall in the following two cases:
  • The analog of Sobolev embedding: $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$ and $q_{0}>q_{1}$. We show that both $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$ and $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$ are finitely strictly singular. See Proposition 4.2.
  • The analog of $L^{p}$-$L^{q}$ embedding on domains: $s_{0}=s_{1}$, $p_{0}\geq p_{1}$ and $q_{0}<q_{1}$. In Proposition 4.4 we show that for the map $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$:
    • When $p_{0}=\infty>p_{1}$, it is finitely strictly singular.
    • When $p_{0}=p_{1}=\infty$ or $p_{0}<\infty$, it is strictly singular but not finitely strictly singular.
We would like to point out that the above results on the “quality” of non-compactness for embeddings between Besov spaces closely align with the findings on strict singularity for embeddings between Lorentz sequence spaces $l^{p,q}$ (see [LN23]). The comparison is both interesting and noteworthy.
It is also important to highlight that the Sobolev embedding $W^{k,p}(\Omega)$ into the space $L^{\frac{np}{n-kp},p}(\Omega)$ is not only the optimal (i.e., smallest) target space within the scale of Lorentz spaces but also the optimal target space among all rearrangement-invariant spaces (such as Lebesgue spaces, Lorentz spaces, Orlicz spaces, and similar spaces). Specifically, this implies that if $Z(\Omega)$ is a rearrangement-invariant space such that $V_{0}^{k,p}(\Omega)\subset Z(\Omega)$, then $L^{\frac{np}{n-kp},p}(\Omega)\subseteq Z(\Omega)$ (see [KP06, Theorem A]). In contrast, the Sobolev embedding into $L^{\frac{np}{n-kp}}(\Omega)$ is optimal within the narrower scale of Lebesgue spaces, and the “quality” of non-compactness differs significantly (i.e., non-strictly singular versus finitely strictly singular). This observation suggests that non-strictly singular embeddings may be associated with the “most optimal” embeddings. This insight could help identify cases where we might expect to find the “most optimal” embeddings between Besov spaces, as seen in the case v in Theorem 1.1 and case iii in Theorem 1.2, comparable with cases in which there is a chance for more optimal target spaces, i.e. when embedding is finitely strictly singular or strictly singular.
It is also worth noting that the techniques used in [LM23], [LM19] and [BG89] for Sobolev embeddings are fundamentally different from those employed in this paper. In particular, the results for Sobolev embeddings $W^{k,p}\to L^{\frac{np}{n-kp}}$ require quite sophisticated methods. This contrasts with the techniques used for Besov embeddings and highlights the advantages of working with Besov spaces compared to Sobolev spaces and Lorentz spaces. This is demonstrated in our forthcoming paper [CLY25], in which, by using the results obtained for Besov spaces and their relation with Sobolev spaces, we improved and extend findings from [BG89] and [LM23].
Additionally, it is noteworthy that, to the best of our knowledge, there is no natural Sobolev embedding that is non-compact, strictly singular, yet not finitely strictly singular. This specific scenario, however, can be naturally observed within the framework of Besov spaces, as demonstrated in Theorem 1.1 and Theorem 1.2, underscoring the structural richness of Besov spaces.
The paper is structured as follows. In the next section, we recall essential definitions and notations used throughout the paper, along with results concerning the wavelet decomposition of Besov spaces and their implications. Section 3 focuses on the study of embeddings between sequence spaces. In Section 4, we present the proofs of the main theorems, Theorems 1.1 and 1.2. Section 5 extends Theorem 1.2 to the setting of homogeneous Besov spaces, resulting in Theorem 5.1. Finally, Section 6 concludes the paper with final remarks and observations.
2. Preliminaries
In this section, we recall some notations, definitions, and auxiliary results.
Throughout the paper, by linear spaces, we mean the real linear spaces. We focus on quasi-Banach spaces as most of the results we used on Banach spaces remain the same for quasi-Banach spaces.
[…]
Definition 2.1.
Let $T:X\to Y$ be a bounded linear map between two (real) quasi-normed spaces $X$ and $Y$.
We say $T$ is strictly singular if there is no infinite-dimensional subspace $Z$ of $X$ such that the restriction $T|_{Z}$ is a topological isomorphism of $Z$ onto $T(Z)$. Equivalently, we can say that
$\inf_{x\in Z;\|x\|_{X}=1}\|T(x)\|_{Y}=0,\quad\text{for every subspaces }Z\subseteq X\text{ such that }\dim Z=\infty.$
We say $T$ is finitely strictly singular if for any $\varepsilon>0$, there exists $N(\varepsilon)\in\mathbb{Z}_{+}$ such that if $E$ is a subspace of $X$ with $\dim E\geq N(\varepsilon)$, then there exists $x\in E$ with $\|x\|_{X}=1$, such that $\|Tx\|_{Y}\leq\varepsilon$.
The finitely strictly singularity can also be expressed in terms of the Bernstein numbers of $T$. Observe that $T$ is finitely strictly singular if and only if $\lim_{n\to\infty}b_{n}(T)=0$.
Definition 2.2.
For $n\geq 1$, the $n$-th Bernstein number $b_{n}(T)$ of a bounded linear map $T:X\to Y$ is
$b_{n}(T):=\sup\big\{\inf_{x\in Z_{n};\|x\|_{X}=1}\left\|T(x)\right\|_{Y}:Z_{n}\subseteq X\text{ is a subspace such that }\dim Z_{n}\geq n\big\}.$
Obviously, we have:
[…]
To classify embedding between Besov spaces, we reduce the discussion to sequence spaces.
Convention 2.3.
Let $q\in(0,\infty]$ and $S$ be a set, we denote by $\ell^{q}(S)$ the spaces $q$-summable sequences $x=(x_{j})_{j\in S}$ with quasinorm $\|x\|_{\ell^{q}(S)}:=(\sum_{j\in S}|x_{j}|^{q})^{1/q}$ when $q<\infty$ and $\|x\|_{\ell^{\infty}(S)}=\sup_{j\in S}|x_{j}|$. When $S=\mathbb{N}_{0}=\{0,1,2,\dots\}$ we use abbreviation $\ell^{q}=\ell^{q}(\mathbb{N}_{0})$.
Definition 2.4.
Let $(X_{j})_{j=0}^{\infty}$ be a family of quasi-normed spaces and let $q\in(0,\infty]$. We denote by $\ell^{q}(X_{j})_{j=0}^{\infty}=\ell^{q}(\mathbb{N}_{0};(X_{j})_{j=0}^{\infty})$ the space of vector-valued sequence $x=(x_{0},x_{1},x_{2},\dots)$ where $x_{j}\in X_{j}$ for each $j\geq 0$, such that
$\|x\|_{\ell^{q}(X_{j})_{j=0}^{\infty}}:=\Big(\sum_{j=0}^{\infty}\|x_{j}\|_{X_{j}}^{q}\Big)^{1/q}.$
Here when $q=\infty$ we replace the sum by supremum over $j\geq 0$.
We also write $x=\sum_{j=0}^{\infty}x_{j}\otimes e_{j}$ where $e_{j}=(0,\dots,0,\underset{j}{1},0,\dots)$ is the standard basis in $\ell^{q}$.
When $(X_{j})_{j=0}^{\infty}$ are all identical for all $j\geq 0$, say $X_{j}\equiv Y$. We use the notation $\ell^{q}(Y)=\ell^{q}(X_{j})_{j=0}^{\infty}$.
Remark 2.5.
Using the notation in the definition, for the double sequence space $\ell^{q}(\ell^{p})$, by the element $e_{j}\otimes e_{k}$ we mean $e_{j}\in\ell^{p}$ and $e_{k}\in\ell^{q}$. More precisely $e_{j}\otimes e_{k}=(0,\dots,0,\underset{k}{e_{j}},0,\dots)$.
Remark 2.6.
Notice that $\ell^{q}(X_{j})_{j=0}^{\infty}$ may not be quasi-normed since $(X_{j})_{j=0}^{\infty}$ may not have uniform triangle inequality. That is, one may not find a $C>0$ such that $\|x_{j}^{\prime}+x_{j}^{\prime\prime}\|_{X_{j}}\leq C(\|x_{j}^{\prime}\|_{X_{j}}+\|x^{\prime\prime}_{j}\|_{X_{j}})$ for all $j\geq 0$ and $x_{j}^{\prime},x_{j}^{\prime\prime}\in X_{j}$. For example, we can take $X_{j}=\ell^{1/j}$ for $j\geq 0$. However, in our case $(X_{j})_{j=0}^{\infty}$ are always subspaces of $\ell^{p}$ for some $p>0$, where we have $\|x_{j}^{\prime}+x_{j}^{\prime\prime}\|_{X_{j}}\leq 2^{\max(1/p-1,0)}(\|x_{j}^{\prime}\|_{X_{j}}+\|x^{\prime\prime}_{j}\|_{X_{j}})$ for all $j$ and $x_{j}^{\prime},x_{j}^{\prime\prime}$.
Convention 2.7.
Let $(X,\|\cdot\|_{X})$ be a quasi-normed space and let $\lambda>0$. We denote by $\lambda X$ the quasi-normed space $(X,\|\cdot\|_{\lambda X})$ where $\|x\|_{\lambda X}:=\lambda\|x\|_{X}$ for every $x\in X$.
Remark 2.8.
For every finite set $S$ and $0<r\leq\infty$, the space $(\#S)^{-1/r}\cdot\ell^{r}(S)$ equals to $L^{r}(S,\mu_{S})$ where $\mu_{S}\{a\}=(\#S)^{-1}$ for every $a\in S$.
Since $(S,\mu_{S})$ is a probability space, for every $0<r_{1}\leq r_{0}\leq\infty$, the embedding
$L^{r_{0}}(S,\mu_{S})=(\#S)^{-1/r_{0}}\cdot\ell^{r_{0}}(S)\hookrightarrow L^{r_{1}}(S,\mu_{S})=(\#S)^{-1/r_{1}}\cdot\ell^{r_{1}}(S)$
has operator norm 1.
Definition 2.9.
Let $s\in\mathbb{R}$ and $p,q\in(0,\infty]$. We denote by $\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ the Besov space which contains all tempered distributions $f\in\mathscr{S}^{\prime}(\mathbb{R}^{n})$ for which
$\|f\|_{\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})}=\|f\|_{\mathscr{B}_{pq}^{s}(\mathbb{R}^{n};\varphi)}:=\Big(\sum_{j=0}^{\infty}\|2^{js}\varphi_{j}\ast f\|_{L^{p}(\mathbb{R}^{n})}^{q}\Big)^{\frac{1}{q}}<\infty.$
Here $(\varphi_{j})_{j=0}^{\infty}\subset\mathscr{S}(\mathbb{R}^{n})$ is a fixed family of Schwartz functions whose Fourier transform $\hat{\varphi}_{j}(\xi)=\int\varphi_{j}(x)e^{-2\pi ix\xi}dx$ satisfy
  • $\operatorname{supp}\hat{\varphi}_{0}\subset B(0,2)$ and $\hat{\varphi}_{0}\equiv 1$ in a neighborhood of $\overline{B(0,1)}$.
  • $\hat{\varphi}_{j}(\xi)=\hat{\varphi}_{0}(2^{-j}\xi)-\hat{\varphi}_{0}(2^{1-j}\xi)$ for $j\geq 1$.
Different choices of $(\varphi_{j})_{j}$ result in different but equivalent norms (see e.g. [Tri10, Proposition 2.3.2]), thus, the space is independent of $(\varphi_{j})_{j}$ even though its norm is dependent on the choice of functions. In this paper, we do not use the exact norm definition from above but the discretized version via wavelet, see Proposition 2.13.
Definition 2.10 (See also [Tri06, Definition 1.95]).
For an open subset $\Omega\subseteq\mathbb{R}^{n}$, we define $\mathscr{B}_{pq}^{s}(\Omega)$ by the restriction of $\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ functions to $\Omega$:
$\mathscr{B}_{pq}^{s}(\Omega):=\{\tilde{f}|_{\Omega}:\tilde{f}\in\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})\},\quad\|f\|_{\mathscr{B}_{pq}^{s}(\Omega)}:=\inf_{\tilde{f}|_{\Omega}=f}\|\tilde{f}\|_{\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})}.$
We define the closed subspace $\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega})\subseteq\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ by
$\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega}):=\{\tilde{f}\in\mathscr{B}_{pq}^{s}(\mathbb{R}^{n}):\operatorname{supp}f\subseteq\overline{\Omega}\}.$
Remark 2.11.
From the above definition, we have the following quotient characterization: for $\Omega\subset\mathbb{R}^{n}$,
(2.1) $\mathscr{B}_{pq}^{s}(\Omega)=\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})/\{\tilde{f}\in\mathscr{B}_{pq}^{s}(\mathbb{R}^{n}):\tilde{f}|_{\Omega}=0\}.$
When $\Omega=\mathbb{R}^{n}$ we have by definition the coincidence of notations
$\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})=\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\mathbb{R}^{n}}).$
Lemma 2.12.
Let $\Omega_{0},\Omega_{1}\subseteq\mathbb{R}^{n}$ be two open subsets such that $\Omega_{0}$ is precompact in $\Omega_{1}$. We have the natural topological embedding $\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega}_{0})\hookrightarrow\mathscr{B}_{pq}^{s}(\Omega_{1})$.
Proof.
An $f\in\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega}_{0})$ is an element in $\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$, thus by definition and (2.1) $\|f\|_{\mathscr{B}_{pq}^{s}(\Omega_{1})}\leq\|f\|_{\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})}$.
On the other hand, take a $\chi\in C_{c}^{\infty}(\Omega_{1})$ such that $\chi|_{\Omega_{0}}\equiv 1$. Recall from e.g. [Tri10, Theorem 2.8.2] that $\tilde{f}\mapsto\chi\tilde{f}]:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})\to\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ is a bounded map whose kernel contains $\{\tilde{f}:\tilde{f}|_{\Omega_{1}}=0\}$. By the quotient characterization (2.1) it descends to a bounded linear map $f\mapsto\chi f]:\mathscr{B}_{pq}^{s}(\Omega_{1})\to\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$, i.e. there is a $C>0$ such that $\|\chi f\|_{\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})}\leq C\|f\|_{\mathscr{B}_{pq}^{s}(\Omega_{1})}$ for all $f\in\mathscr{B}_{pq}^{s}(\Omega_{1})$.
Since $f=\chi f$ for every $f\in\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega}_{0})$, this is to say $\|f\|_{\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})}=\|\chi f\|_{\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})}\leq C\|f\|_{\mathscr{B}_{pq}^{s}(\Omega_{1})}$ for all $f\in\widetilde{\mathscr{B}}_{pq}^{s}(\overline{\Omega}_{0})$, completing the proof. ∎
The mapping between Besov spaces can be reduced to the corresponding sequence spaces via wavelet decomposition theorems. The following theorem is due to Ciesielski and Figiel (see [CF83a, CF83b]).
In the following we use indices sets $Q_{j}$ and $G_{j}$ for $j\geq 0$, given by
(2.2) $Q_{j}:=2^{-j}\mathbb{Z}^{n}\text{ for }j\geq 0;\qquad G_{0}:=\{0,1\}^{n},\quad G_{j}:=\{0,1\}^{n}\backslash\{0\}^{n}\text{ for }j\geq 1.$
Proposition 2.13 ([Tri08, Theorem 1.20]).
Let $\varepsilon>0$. There is a set $\{\psi_{jm}^{g}:j\geq 0,m\in Q_{j},g\in G_{j}\}\subset C_{c}^{\infty}(\mathbb{R}^{n})$ (depending on $\varepsilon$) which forms a (real) orthonormal basis to $L^{2}(\mathbb{R}^{n})$ satisfying the following.
  1. (i)
    $\psi_{jm}^{g}(x)=2^{jn/2}\psi_{00}^{g}(2^{j}(x-m))$ for all $j\geq 0$, $m\in Q_{j}$ and $g\in G_{j}$.
  2. (ii)
    There is a $L\geq 1$ such that $\operatorname{supp}\psi_{00}^{g}\subset B(0,2^{L})$ for all $g\in G_{0}$. As a result, $\operatorname{supp}\psi_{jm}^{g}\subset B(m,2^{L-j})$.
  3. (iii)
    For every $p,q\in[\varepsilon,\infty],s\in[-\varepsilon^{-1},\varepsilon^{-1}]$ and $f\in\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$, we have expansion (which is orthonormal in $L^{2}$)
    $f=\sum_{j=0}^{\infty}\sum_{m\in Q_{j},g\in G_{j}}(f,\psi_{jm}^{g})\cdot\psi_{jm}^{g},$
    where the series converges in the sense of tempered distributions.
  4. (iv)
    For every $p,q\in[\varepsilon,\infty],s\in[-\varepsilon^{-1},\varepsilon^{-1}]$,
    (2.3) $f\in\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})\text{ if and only if }\bigg(\sum_{j=0}^{\infty}2^{j(s-\frac{n}{p}+\frac{n}{2})q}\Big|\sum_{m\in Q_{j},g\in G_{j}}|(f,\psi_{jm}^{g})|^{p}\Big|^{q/p}\bigg)^{1/q}<\infty.$
    In other words we obtain the following isomorphisms
    (2.4) $\begin{array}[]{rcl}\Lambda=\Lambda_{\varepsilon}:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})&\xrightarrow{\simeq}&\ell^{q}\big(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(Q_{j}\times G_{j})\big)_{j=0}^{\infty};\\ f&\longmapsto&\{(f,\psi_{jm}^{g})_{L^{2}}:m\in Q_{j},\ g\in G_{j}\}_{j=0}^{\infty};\\ \sum_{j=0}^{\infty}\sum_{m,g}a_{jm}^{g}\psi_{jm}^{g}&\mathrel{\reflectbox{$\longmapsto$}}&\{a_{jm}^{g}:m\in Q_{j},\ g\in G_{j}\}_{j=0}^{\infty}.\end{array}$
    In particular, by fixing identifications $Q_{j}\times G_{j}\simeq\mathbb{Z}$ for each $j\geq 0$ we obtain the isomorphisms $\Lambda:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})\xrightarrow{\simeq}\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(\mathbb{Z}))_{j=0}^{\infty}$.
Remark 2.14.
The right hand side of (2.3) are actually the so-called Besov sequence norm. See [Tri08, Definition 1.18] for details.
For more information about construction and structure of $\psi_{jm}^{g}$ and $G_{j}$ we directions readers to [Tri08, Section 1.2.1].
In the next we sometimes restrict the $\Lambda$ or $\Lambda^{-1}$ from (2.4) to certain subspaces. Here we use the existence of the extension operator on Lipschitz domains.
Lemma 2.15 ([Ryc99, Theorem 2.2]).
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain and let $U\supset\overline{\Omega}$ be an open neighborhood. For every $\varepsilon>0$, there is an extension operator $E$ such that $E:\mathscr{B}_{pq}^{s}(\Omega)\to\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})$ is defined and bounded for all $p,q\in[\varepsilon,\infty]$ and $s\in[-\varepsilon^{-1},\varepsilon^{-1}]$.
In fact, we allow $\varepsilon=0$ as we can build a universal extension operator, see [Ryc99, Theorem 4.1].
Here, we also remark that $E:\mathscr{B}_{pq}^{s}(\Omega)\to\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})$ is always a topological embedding since it is the right inverse to the restriction mapping $\tilde{f}\mapsto\tilde{f}|_{\Omega}]:\mathscr{B}_{pq}^{s}(\mathbb{R}^{n})\twoheadrightarrow\mathscr{B}_{pq}^{s}(\Omega)$.
Proposition 2.16.
Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain, and $U$ and $V$ be bounded open sets such that $V\Subset\Omega\Subset U$. Let $\varepsilon>0$, $\Lambda$ be as in (2.4) (depending on $\varepsilon$), and let $E$ be the extension operator as in Lemma 2.15 (depending on $\Omega,U$).
Then there are an $A\geq 1$ and sets $R_{j}\subset S_{j}\subset Q_{j}\times G_{j}$ for $j\geq 0$, such that, for every $p,q\in[\varepsilon,\infty]$ and $s\in[-\varepsilon^{-1},\varepsilon^{-1}]$:
  1. (i)
    $R_{j}=2^{(j-A)n}$ for $j\geq A$ and we have $\Lambda^{-1}:\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(R_{j}))_{j=A}^{\infty}\to\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})$. As a result $\Lambda^{-1}:\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(R_{j}))_{j=A}^{\infty}\to\mathscr{B}_{pq}^{s}(\Omega)$ is a topological embedding.
  2. (ii)
    $S_{j}=2^{(j+A)n}$ for $j\geq 0$ and we have $\Lambda:\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})\to\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(S_{j}))_{j=0}^{\infty}$. As a result $\Lambda\circ E:\mathscr{B}_{pq}^{s}(\Omega)\to\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(S_{j}))_{j=0}^{\infty}$ is a topological embedding.
Here, the constant $A$ depends on the number $L$ in Proposition 2.13 ii, the inner radius of $V$ and the diameter of $U$.
Remark 2.17.
It is possible to show that for a bounded Lipschitz domain $\Omega$, there is a linear map $\tilde{\Lambda}$ such that $\tilde{\Lambda}:\mathscr{B}_{pq}^{s}(\Omega)\xrightarrow{\simeq}\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=0}^{\infty}$ are isomorphisms for all $p,q\in[\varepsilon,\infty]$ and $s\in[-\varepsilon^{-1},\varepsilon^{-1}]$. However its proof may be complicated as it requires some re-ordering of certain subsets of $S_{j}$ (which contains $R_{j}$).
But for our needs, the weaker results i and ii are satisfactory.
Proof of Proposition 2.16.
Let $L\geq 1$ be from Proposition 2.13 ii. We define $\mathfrak{R}_{j},\mathfrak{S}_{j}\subset Q_{j}$ by $\mathfrak{R}_{j}:=\{m\in Q_{j}:B(m,2^{L-j})\subset V\}$ and $\mathfrak{S}_{j}:=\{m\in Q_{j}:B(m,2^{L-j})\cap U\neq\varnothing\}$.
Recall $Q_{j}=2^{-j}\mathbb{Z}^{n}$, thus for a large enough $A\geq 1$ (depending on $U,V,L$) we have $\mathfrak{R}_{j}\geq 2^{(j-A)n}$ (for $j\geq A$) and $\mathfrak{S}_{j}\leq 2^{(j+A-1)n}$ (for $j\geq 0$).
By Proposition 2.13 ii we have $\operatorname{supp}\psi_{jm}^{g}\subset V$ for all $m\in\mathfrak{R}_{j}$ and $g\in G_{j}$, thus by (2.3) $\sum_{j=A}^{\infty}\sum_{m\in\mathfrak{R}_{j},g\in G_{j}}a_{jm}^{g}\psi_{jm}^{g}\in\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})$ holds for all $\{a_{jm}^{g}\}_{j=A}^{\infty}\in\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(\mathfrak{R}_{j}\times G_{j}))_{j=0}^{\infty}$. Taking $R_{j}\subseteq\mathfrak{R}_{j}\times G_{j}$ be any set such that $R_{j}=2^{(j-A)n}$ and applying Proposition 2.13 iv we get $\Lambda^{-1}:\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(R_{j}))_{j=A}^{\infty}\to\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})$. Applying Lemma 2.12 we have topological embedding $\widetilde{\mathscr{B}}_{pq}^{s}(\overline{V})\hookrightarrow\mathscr{B}_{pq}^{s}(\Omega)$. This completes the proof of i.
Similarly by Proposition 2.13 ii and iii, for every $f\in\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})$, we have $(f,\phi_{jm}^{g})_{L^{2}}=0$ if $m\notin\mathfrak{S}_{j}$, hence by Proposition 2.13 iii $f=\sum_{j=0}^{\infty}\sum_{m\in\mathfrak{S}_{j},g\in G_{j}}(f,\psi_{jm}^{g})_{L^{2}}\cdot\psi_{jm}^{g}$. Taking $S_{j}\supseteq\mathfrak{S}_{j}\times G_{j}$ be any set in $Q_{j}\times G_{j}$ such that $S_{j}=2^{(j+A)n}$ and applying Proposition 2.13 iv we get $\Lambda:\widetilde{\mathscr{B}}_{pq}^{s}(\overline{U})\to\ell^{q}(2^{j(s-\frac{n}{p}+\frac{n}{2})}\cdot\ell^{p}(S_{j}))_{j=0}^{\infty}$. Applying Lemma 2.15 which gives $E:\mathscr{B}_{pq}^{s}(\Omega)\to\mathscr{B}_{pq}^{s}(\overline{U})$, we get the required boundedness of $\Lambda\circ E$. Since $\Lambda$ and $E$ are both topological embeddings, so does $\Lambda\circ E$. This completes the proof of ii. ∎
Now we reduce the discussion from function spaces to sequence spaces.
The following well-known result can also be deduced from the wavelet decomposition.
Corollary 2.18.
Let $s_{0},s_{1}\in\mathbb{R}$ and $0<p_{0},p_{1},q_{0},q_{1}\leq\infty$.
  1. (i)
    The inclusion $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\subseteq\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$ holds if and only if
    (2.5) $\big(s_{0}-s_{1}=\tfrac{n}{p_{0}}-\tfrac{n}{p_{1}}\geq 0\text{ and }q_{0}\leq q_{1}\big)\quad\text{or}\quad\big(s_{0}-s_{1}>\tfrac{n}{p_{0}}-\tfrac{n}{p_{1}}\geq 0\big).$
    Note that both cases are non-compact.
  2. (ii)
    Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain. Then $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\subseteq\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$ holds if and only if
    (2.6) $s_{0}-s_{1}=\max(0,\tfrac{n}{p_{0}}-\tfrac{n}{p_{1}})\text{ and }q_{0}\leq q_{1};$
    (2.7) $\text{or}\quad s_{0}-s_{1}>\max(0,\tfrac{n}{p_{0}}-\tfrac{n}{p_{1}}).$
    Moreover the embedding is non-compact if and only if (2.6) holds.
Proof.
The proof is standard, see e.g. [Tri10, Chapter 2.7.1], [Tri06, Chapter 1.11.1] or [Tri20, Remark 2.87].
Indeed one can check that $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}}+\frac{n}{2})}\cdot\ell^{p_{0}})_{j=0}^{\infty}\subset\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}}+\frac{n}{2})}\cdot\ell^{p_{1}})_{j=0}^{\infty}$ holds if and only if either (2.5) holds; and the map is always non-compact. By Proposition 2.13 iv we get i.
Similarly $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}}+\frac{n}{2})}\cdot\ell^{p_{0}}\{1,\dots,2^{nj}\})_{j=0}^{\infty}\subset\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}}+\frac{n}{2})}\cdot\ell^{p_{1}}\{1,\dots,2^{nj}\})_{j=0}^{\infty}$ if and only if either (2.6) or (2.7) holds; and the map is compact if and only if (2.7) holds.
By Proposition 2.16 i we see that $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\subset\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$ implies (2.6) or (2.7). By Proposition 2.16 ii and using $\Lambda\circ E$ we see that (2.6) or (2.7) implies $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\subset\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$. The compactness of the embedding $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$ follows the same argument via Proposition 2.16 i and ii. ∎
3. Embeddings Between Sequence Spaces
From the wavelet decomposition in the previous section, the problems are reduced to the embedding between sequence spaces. The goal of this section is to prove the following: let $0<q_{0}<q_{1}\leq\infty$.
  • (Proposition 3.5) Let $p_{0}<p_{1}$. Then $\ell^{q_{0}}(\ell^{p_{0}})\hookrightarrow\ell^{q_{1}}(\ell^{p_{1}})$ is finitely strictly singular.
  • (Corollary 3.4) Assume $p_{1}<\infty$. Then $\ell^{q_{0}}(L^{\infty}[0,1])\hookrightarrow\ell^{q_{1}}(L^{p_{1}}[0,1])$ is finitely strictly singular.
  • (Propositions 3.7 and 3.8) Assume either $p_{1}\leq p_{0}<\infty$ or $p_{0}=p_{1}=\infty$. Then,
    $\ell^{q_{0}}(2^{-jn/p_{0}}\cdot\ell^{p_{0}}\{1,\dots,2^{j}\})_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{-jn/p_{1}}\cdot\ell^{p_{1}}\{1,\dots,2^{j}\})_{j=0}^{\infty}$
    is strictly singular but not finitely strictly singular.
Note that when $q_{0}=q_{1}$ all the above embeddings are not strictly singular.
These results are inspired from Plichko [Pli04] and Lefèvre, Rodríguez-Piazza [LRP14]. Although some of the techniques might not be new, to the best of the authors’ knowledge, the proof is not found in the standard literature. Thus we attach a full proof in detail.
Lemma 3.1.
Let $(T_{j}:X_{j}\to Y_{j})_{j=0}^{\infty}$ be a family of bounded linear maps between quasi-normed spaces such that $0<\limsup_{j\to\infty}\|T_{j}\|_{X_{j}\to Y_{j}}<\infty$. Define $\widehat{T}:(x_{j})_{j=0}^{\infty}\mapsto(T_{j}x_{j})_{j=0}^{\infty}$.
Then for every $0<q\leq\infty$, $\widehat{T}:\ell^{q}(X_{j})_{j=0}^{\infty}\to\ell^{q}(Y_{j})_{j=0}^{\infty}$ is not strictly singular.
Proof.
Let $e_{j}=(0,\dots,0,\underset{j}{1},0,\dots)\in\ell^{\infty}$ be the standard elements. By assumption, there exist $c>0$ and $x_{j}\in X_{j}$ such that for all $j\geq 0$, $\|x_{j}\|_{X_{j}}=1$ and $\|T_{j}x_{j}\|_{Y_{j}}\geq c$.
Next, consider the subspace $V:=\operatorname{span}\{x_{j}\otimes e_{j}:j\geq 0\}\subset\ell^{q}(X_{j})_{j=0}^{\infty}$. For any sequence $(a_{j})_{j=0}^{\infty}\subset\mathbb{R}$, we have $\|\sum_{j=0}^{\infty}a_{j}x_{j}\otimes e_{j}\|_{\ell^{q}(X_{j})_{j}}=\|(a_{j})_{j}\|_{\ell^{q}}$ and $\|\widehat{T}\sum_{j=0}^{\infty}a_{j}x_{j}\otimes e_{j}\|_{\ell^{q}(Y_{j})_{j}}\geq c\|(a_{j})_{j}\|_{\ell^{q}}$, giving the non-strictly singularity. ∎
In order to prove the required propositions, need to prove some results on the ambient spaces and the apply them to their subspaces.
In the proof we apply the diagonal theorem from [LRP14, Theorem 3.3]:
Lemma 3.2.
Let $(X_{j},Y_{j})_{j=0}^{\infty}$ be Banach spaces. Let $T_{j}:X_{j}\to Y_{j}$ ($j\geq 0$) be bounded linear maps where the operator norms are uniformly bounded over $j\geq 0$, i.e. $\sup_{j\geq 0}\sup_{\|x_{j}\|_{X_{j}}=1}\|T_{j}x_{j}\|_{Y_{j}}<\infty$. For given $0<p<q\leq\infty$ we define $\widehat{T}_{p,q}:\ell^{p}(X_{j})_{j=0}^{\infty}\to\ell^{q}(X_{j})_{j=0}^{\infty}$ by $\widehat{T}_{p,q}:(x_{j})_{j=0}^{\infty}\mapsto(T_{j}x_{j})_{j=0}^{\infty}$.
Then $\widehat{T}_{p,q}$ is finitely strictly singular if $(T_{j})_{j}$ are uniformly finitely strictly singular.
More precisely, that is $\lim_{n\to\infty}b_{n}(\widehat{T}_{p,q})=0$ if $\lim_{n\to\infty}\sup_{j\geq 0}b_{n}(T_{j})=0$.
Remark 3.3.
When $1\leq p<q\leq\infty$, this is exactly the result in [LRP14]. It is possible that the diagonal theorem holds for $(X_{j},Y_{j})_{j=0}^{\infty}$ as well. But for our application this is already enough.
Proof of Lemma 3.2.
We prove the result by assuming [LRP14, Theorem 3.3].
Using composition $\ell^{p}(X_{j})_{j=0}^{\infty}\hookrightarrow\ell^{\max(p,1)}(X_{j})_{j=0}^{\infty}\xrightarrow{\widehat{T}_{\max(p,1),\infty}}\ell^{\infty}(Y_{j})_{j=0}^{\infty}$ and applying [LRP14, Theorem 3.3] on $\widehat{T}_{\max(p,1),\infty}$ we see that $\widehat{T}_{p,\infty}$ is finitely strictly singular for all $0<p<\infty$.
When $q<\infty$, without loss of generality we can assume $\|T_{j}\|_{X_{j}\to Y_{j}}\leq 1$ for all $j\geq 0$. Therefore
$\|\widehat{T}_{p,q}x\|_{\ell^{q}(Y_{j})_{j=0}^{\infty}}^{q}=$ $\sum_{j=0}^{\infty}\|T_{j}x_{j}\|_{Y_{j}}^{q}\leq\sum_{j=0}^{\infty}\|x_{j}\|_{X_{j}}^{p}\|T_{j}x_{j}\|_{Y_{j}}^{q-p}\leq\|x\|_{\ell^{p}(X_{j})_{j=0}^{\infty}}^{p}\sup_{j\geq 0}\|T_{j}x_{j}\|_{Y_{j}}^{q-p}$
$=$ $\|x\|_{\ell^{p}(X_{j})_{j=0}^{\infty}}^{p}\|\widehat{T}_{p,\infty}x\|_{\ell^{\infty}(Y_{j})_{j=0}^{\infty}}^{q-p}.$
By assumption, for every $n$-dimensional subspace $V\subset\ell^{p}(X_{j})_{j=0}^{\infty}$ there is a unit vector $x\in V$ such that $\|\widehat{T}_{p,\infty}x\|_{\ell^{\infty}(Y_{j})_{j}}\leq b_{n}(\widehat{T}_{p,\infty})\|x\|_{\ell^{p}(X_{j})_{j}}$. Therefore $\|\widehat{T}_{p,q}x\|_{\ell^{q}(Y_{j})_{j}}\leq\|\widehat{T}_{p,\infty}x\|_{\ell^{\infty}(Y_{j})_{j}}^{1-p/q}\leq b_{n}(\widehat{T}_{p,\infty})^{1-p/q}$. Since $\widehat{T}_{p,\infty}$ is finitely strictly singular by the above, the right hand side tends to 0, which concludes the proof. ∎
Corollary 3.4.
For every $0<p<\infty$ and $0<q_{0}<q_{1}\leq\infty$, $\ell^{q_{0}}(L^{\infty}[0,1])\hookrightarrow\ell^{q_{1}}(L^{p}[0,1])$ is finitely strictly singular.
Proof.
It is known that $L^{\infty}[0,1]\hookrightarrow L^{p}[0,1]$ is finitely strictly singular. See e.g. the proof [Rud91, Theorem 5.2]. Although the statement [Rud91, Theorem 5.2] merely says that $L^{\infty}[0,1]\hookrightarrow L^{p}[0,1]$ is strictly singular, its proof gives the control $b_{n}(L^{\infty}[0,1]\hookrightarrow L^{p}[0,1])\leq n^{-1/\max(p,2)}$, which tends to zero as $n\to\infty$. See also [HST17, Corollary 29].
Applying Lemma 3.2 the map $\ell^{q_{0}}(L^{\infty}[0,1])\hookrightarrow\ell^{q_{1}}(L^{\max(p,2)}[0,1])$ is also finitely strictly singular. Therefore by composing the map $\ell^{q_{1}}(L^{\max(p,2)}[0,1])\hookrightarrow\ell^{q_{1}}(L^{p}[0,1])$ we conclude the proof. ∎
The proof of Theorem 1.2 ii is reduced to the following statement on the sequence spaces.
Proposition 3.5.
Let $0<p_{0}<p_{1}\leq\infty$ and $0<q_{0}<q_{1}\leq\infty$.
$b_{n}\big(\ell^{q_{0}}(\ell^{p_{0}})\hookrightarrow\ell^{q_{1}}(\ell^{p_{1}})\big)\leq n^{-\min(\frac{1}{p_{0}},\frac{1}{q_{0}})(1-\max(\frac{q_{0}}{q_{1}},\frac{p_{0}}{p_{1}}))}.$
In particular $\ell^{q_{0}}(\ell^{p_{0}})\hookrightarrow\ell^{q_{1}}(\ell^{p_{1}})$ is finitely strictly singular.
Proof.
The Hölder’s inequality yields $\|x\|_{\ell^{q/\theta}(\ell^{p/\theta})}\leq\|x\|_{\ell^{q}(\ell^{p})}^{\theta}\|x\|_{\ell^{\infty}}^{1-\theta}$ for all $0<p,q\leq\infty$ and $0<\theta<1$. Therefore,
(3.1) $\|x\|_{\ell^{q_{1}}(\ell^{p_{1}})}\leq\|x\|_{\ell^{q_{0}\cdot\min(\frac{q_{1}}{q_{0}},\frac{p_{1}}{p_{0}})}(\ell^{p_{0}\cdot\min(\frac{q_{1}}{q_{0}},\frac{p_{1}}{p_{0}})})}\leq\|x\|_{\ell^{q_{0}}(\ell^{p_{0}})}^{\max(\frac{q_{0}}{q_{1}},\frac{p_{0}}{p_{1}})}\|x\|_{c_{0}}^{1-\max(\frac{q_{0}}{q_{1}},\frac{p_{0}}{p_{1}})}.$
Recall from e.g. [Pli04, Lemma 4], that if $V\subset c_{0}(\mathbb{N}_{0})$ is a linear subspace of dimension $\geq n$, then there is a $x=(x_{j})_{j=0}^{\infty}\in V$ such that $\max_{j\geq 0}|x_{j}|$ is attained in at least $n$ points.
Let $V\subset\ell^{q_{0}}(\ell^{p_{0}})$ be a given $n$ dimensional subspace. Therefore, there is a nonzero $x=(x_{j,k})_{j,k=0}^{\infty}$ such that $\max_{j,k\geq 0}|x_{j,k}|$ is attained in at least $n$ places. Suppose $\|x\|_{\ell^{q_{0}}(\ell^{p_{0}})}=1$, then we have
$1=\|x\|_{\ell^{q_{0}}(\ell^{p_{0}})}\geq\|x\|_{\ell^{\max(p_{0},q_{0})}}\geq(n\|x\|_{c_{0}}^{\max(p_{0},q_{0})})^{\max(p_{0},q_{0})^{-1}}\quad\Rightarrow\quad\|x\|_{c_{0}}\leq n^{-\max(p_{0},q_{0})^{-1}}.$
Using (3.1) we conclude that
$b_{n}\big(\ell^{q_{0}}(\ell^{p_{0}})\hookrightarrow\ell^{q_{1}}(\ell^{p_{1}})\big)\leq n^{-\max(p_{0},q_{0})^{-1}(1-\max(\frac{q_{0}}{q_{1}},\frac{p_{0}}{p_{1}}))}.$
This completes the proof. ∎
Corollary 3.6.
Let $s\geq t$, $0<p_{0}<p_{1}\leq\infty$ and $0<q_{0},q_{1}\leq\infty$ such that either $s>t$ or $q_{0}<q_{1}$. Then the following embedding is finitely strictly singular, in particular strictly singular:
$\ell^{q_{0}}(2^{js}\cdot\ell^{p_{0}})_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{jt}\cdot\ell^{p_{1}})_{j=0}^{\infty}.$
Proof.
First we consider the case $q_{0}<q_{1}$ and $s=t$. Notice that the map $J_{s}(x_{j})_{j=0}^{\infty}:=(2^{js}x_{j})_{j=0}^{\infty}$ defines an isomorphism $J_{s}:\ell^{q}(2^{js}\cdot\ell^{p})_{j=0}^{\infty}\to\ell^{q}(\ell^{p})$ for every $0<p,q\leq\infty$. Therefore $\ell^{q_{0}}(2^{js}\cdot\ell^{p_{0}})_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{js}\cdot\ell^{p_{1}})_{j=0}^{\infty}$ and $\ell^{q_{0}}(\ell^{p_{0}})\hookrightarrow\ell^{q_{1}}(\ell^{p_{1}})$ share the same Bernstein numbers. The finitely strictly singularity follows immediately from Proposition 3.5.
When $s>t$, we can decompose $\ell^{q_{0}}(2^{js}\cdot\ell^{p_{0}})_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{jt}\cdot\ell^{p_{1}})_{j=0}^{\infty}$ into
$\ell^{q_{0}}(2^{js}\cdot\ell^{p_{0}})_{j=0}^{\infty}\hookrightarrow\ell^{\min(\frac{q_{1}}{2},1)}(2^{jt}\cdot\ell^{p_{0}})_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{jt}\cdot\ell^{p_{1}})_{j=0}^{\infty}.$
Here $\ell^{q_{0}}(2^{js}\cdot\ell^{p_{0}})_{j=0}^{\infty}\hookrightarrow\ell^{\min(q_{1}/2,1)}(2^{jt}\cdot\ell^{p_{0}})_{j=0}^{\infty}$ is bounded because by Hölder’s inequality:
$\|x\|_{\ell^{\min(q_{1}/2,1)}(2^{jt}\cdot\ell^{p_{0}})_{j=0}^{\infty}}^{\min(q_{1}/2,1)}=$ $\sum_{j=0}^{\infty}(2^{jt}\|x_{j}\|_{\ell^{p_{0}}})^{\min(q_{1}/2,1)}\leq\sup_{j\geq 0}(2^{js}\|x_{j}\|_{\ell^{p_{0}}})^{\min(q_{1}/2,1)}\sum_{k=0}^{\infty}2^{-k(s-t)\cdot\min(q_{1}/2,1)}$
$\lesssim$ $_{s,t,q_{1}}\|x\|_{\ell^{\infty}(2^{js}\cdot\ell^{p_{0}})_{j=0}^{\infty}}^{\min(q_{1}/2,1)}\leq\|x\|_{\ell^{q_{0}}(2^{js}\cdot\ell^{p_{0}})_{j=0}^{\infty}}^{\min(q_{1}/2,1)}.$
The result then follows immediately since $\ell^{\min(\frac{q_{1}}{2},1)}(2^{jt}\cdot\ell^{p_{0}})_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{jt}\cdot\ell^{p_{1}})_{j=0}^{\infty}$ is already known to be finitely strictly singular. ∎
The key to prove Theorem 1.1 iv is the following:
Proposition 3.7.
Let $0<q_{0}<q_{1}\leq\infty$. Let $X$ be a fixed quasi-Banach space and let $X_{\bullet}=(X_{j})_{j=0}^{\infty}\subset X$ be a sequence of finite dimensional subspaces with the norm induced from $X$.
Then the following embedding is strictly singular:
(3.2) $\ell^{q_{0}}(X_{j})_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(X_{j})_{j=0}^{\infty}.$
When $\sup_{j}\dim X_{j}=\infty$ and $q_{1}<\infty$ the map is not finitely strictly singular. For the special case $X_{j}=\ell^{p}\{1,\dots,2^{j}\}$, see Proposition 3.8 below.
Proof.
Let $\varepsilon>0$ and $V\subset\ell^{q_{0}}(X_{j})_{j=0}^{\infty}$ be an infinite dimensional subspace. We need to find a non-zero $x=(x_{j})_{j=0}^{\infty}\in V$ such that $\|x\|_{\ell^{q_{1}}(X)}<\varepsilon\|x\|_{\ell^{q_{0}}(X)}$.
Let $N>1$ be such that $N^{\frac{1}{q_{1}}-\frac{1}{q_{0}}}<\frac{\varepsilon}{2}$. Let $\delta=\delta_{N}\in(0,1)$ be such that
(3.3) $\frac{(1+(N+1)\delta^{\frac{\min(1,q_{1})}{q_{0}}})^{\frac{q_{1}}{\min(1,q_{1})}}}{(1-(N+1)\delta^{\min(1,\frac{1}{q_{0}})})^{\frac{q_{0}}{\min(1,q_{0})}}}<2.$
Such a $\delta$ always exists since if we let $\delta\to 0$ then the left hand side tends to $1$ which is less than $2$.
Since $q_{0}<\infty$ and $X_{j}$ are all finite dimensional, we can find out a sequence $(j_{n})_{n=1}^{N}\subset\mathbb{N}_{0}$ and a $(x^{(n)})_{n=1}^{N}\subset V$ with $\|x^{(n)}\|_{\ell^{q_{0}}(X)}=1$ for each $n\geq 1$, such that
  • $0\leq j_{1}<j_{2}<\dots<j_{N}$;
  • $x^{(n)}_{j}\equiv 0$ for every $0\leq j\leq j_{n-1}$ (here we set $j_{0}=-1$);
  • $\sum_{j=j_{n}+1}^{\infty}\|x^{(n)}_{j}\|_{X_{j}}^{q_{0}}\leq\delta$, or equivalently, $\sum_{j=j_{n-1}+1}^{j_{n}}\|x^{(n)}_{j}\|_{X_{j}}^{q_{0}}\geq 1-\delta$.
The construction is done by induction. We start with a unit vector $x^{(1)}\in V$, since $q_{0}<\infty$ there is a $j_{1}\geq 1$ such that $\sum_{j=j_{1}+1}^{\infty}\|x^{(1)}_{j}\|_{X_{j}}^{q_{0}}\leq\delta$.
Suppose for $n\geq 2$ and we have obtained $x^{(1)},\dots,x^{(n-1)}$ and $j_{1}<\dots<j_{n-1}$. Consider the projection $\mathcal{P}_{n-1}:\ell^{q_{0}}(\mathbb{N}_{0};X)\twoheadrightarrow\ell^{q_{0}}(\{0,\dots,j_{n-1}\};X)$, we have $\operatorname{ker}(\mathcal{P}_{n-1}|_{V})\neq\{0\}$ since $\dim\mathcal{P}_{n-1}(V)=\sum_{j=0}^{j_{n-1}}\dim X_{j}<\infty$ but $\dim V=\infty$. Therefore, there must be a unit vector $x^{(n)}\in V$ such that $\mathcal{P}_{n-1}x^{(n)}=0$, i.e. $x^{(n)}_{j}\equiv 0$ for every $0\leq j\leq j_{n-1}$. Take a $j_{n}>j_{n-1}$ be such that $\sum_{j=j_{n}+1}^{\infty}\|x^{(n)}_{j}\|_{X}^{q_{0}}\leq\delta$ we close the induction.
Now we see that $x:=x^{(1)}+\dots+x^{(N)}\in V$ is nonzero since
$\|x\|_{\ell^{q_{0}}(X)}\geq\|x\|_{\ell^{q_{0}}(\{1,\dots,j_{1}\};X)}=\|x^{(1)}\|_{\ell^{q_{0}}(\{1,\dots,j_{1}\};X)}\geq(1-\delta)^{1/q_{0}}>0.$
We are going to prove that $x$ is what we needed.
Note that for every $a,b,q>0$, we have $(a+b)^{\min(1,q)}\leq a^{\min(1,q)}+b^{\min(1,q)}$. Therefore for each $1\leq n\leq N$,
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}$ $=\Big\|\sum_{k=1}^{n}x^{(k)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}$
$\geq\|x^{(n)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}-\Big\|\sum_{k=1}^{n-1}x^{(k)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}$
$\geq\|x^{(n)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{0})}-\sum_{k=1}^{n-1}\|x^{(k)}\|_{\ell^{q_{0}}(\{j_{k}+1,j_{k}+2,\dots\};X)}^{\min(1,q_{0})}$
$\geq(1-\delta)^{\frac{\min(1,q_{0})}{q_{0}}}-N\delta^{\frac{\min(1,q_{0})}{q_{0}}}\geq 1-(N+1)\delta^{\min(1,\frac{1}{q_{0}})}.$
Taking the sum over $1\leq n\leq N$, we see that
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{0}}(X)}^{q_{0}}\geq\sum_{n=1}^{N}\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{q_{0}}\geq N(1-(N+1)\delta^{\min(1,q_{0}^{-1})})^{\frac{q_{0}}{\min(1,q_{0})}}.$
On the other hand, using the property $\|x\|_{\ell^{q_{1}}(S;X)}\leq\|x\|_{\ell^{q_{0}}(S;X)}$ for every $x\in\ell^{q_{0}}(X)$ and set $S$, we have, for each $1\leq n\leq N$,
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}=$ $\Big\|\sum_{k=1}^{n}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}$
$\leq\|x^{(n)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}$ $+\Big\|\sum_{k=1}^{n-1}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}$
$\leq\|x^{(n)}\Big\|_{\ell^{q_{0}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{\min(1,q_{1})}$ $+\sum_{k=1}^{n-1}\|x^{(k)}\|_{\ell^{q_{1}}(\{j_{k}+1,j_{k}+2,\dots\};X)}^{\min(1,q_{1})}\leq 1+N\delta^{\frac{\min(1,q_{1})}{q_{0}}}.$
Similarly
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{N}+1,j_{N}+2,\dots\};X)}^{\min(1,q_{1})}\leq\sum_{k=1}^{N}\|x^{(k)}\|_{\ell^{q_{1}}(\{j_{N}+1,j_{N}+2,\dots\};X)}^{\min(1,q_{1})}\leq\delta^{\frac{\min(1,q_{1})}{q_{0}}}N.$
Therefore, using the convention $j_{N+1}=\infty$,
$\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{1}}(X)}\leq\bigg(\sum_{n=1}^{N+1}\Big\|\sum_{k=1}^{N}x^{(k)}\Big\|_{\ell^{q_{1}}(\{j_{n-1}+1,\dots,j_{n}\};X)}^{q_{1}}\bigg)^{1/q_{1}}\leq N^{1/q_{1}}(1+(N+1)\delta^{\frac{\min(1,q_{1})}{q_{0}}})^{\frac{1}{\min(1,q_{1})}}.$
Plugging in (3.3) we conclude that
$\frac{\|x^{(1)}+\dots+x^{(N)}\|_{\ell^{q_{1}}(X)}}{\|x^{(1)}+\dots+x^{(N)}\|_{\ell^{q_{0}}(X)}}\leq N^{\frac{1}{q_{1}}-\frac{1}{q_{0}}}\frac{(1+(N+1)\delta^{\frac{\min(1,q_{1})}{q_{0}}})^{\frac{q_{1}}{\min(1,q_{1})}}}{(1-(N+1)\delta^{\min(1,\frac{1}{q_{0}})})^{\frac{q_{0}}{\min(1,q_{0})}}}<\frac{\varepsilon}{2}\cdot 2=\varepsilon.$
Therefore $x=x^{(1)}+\dots+x^{(N)}\in V$ is what we needed to achieve $\|x\|_{\ell^{q_{1}}(\ell^{p})}<\varepsilon\|x\|_{\ell^{q_{0}}(\ell^{p})}$. ∎
The following statement generalizes Proposition 3.7. For convenience, we only prove the case $m_{j}=2^{j}$.
Proposition 3.8.
Let $0<p_{1}\leq p_{0}\leq\infty$ and $0<q_{0}<q_{1}\leq\infty$ such that $(p_{0},p_{1})\notin\{\infty\}\times(0,\infty)$.
Then for every $A\geq 0$, the embedding is not finitely strictly singular:
(3.4) $\ell^{q_{0}}(2^{-\frac{j}{p_{0}}}\cdot\ell^{p_{0}}\{1,\dots,2^{j}\})_{j=A}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{-\frac{j}{p_{1}}}\cdot\ell^{p_{1}}\{1,\dots,2^{j}\})_{j=A}^{\infty}.$
Proof.
When $p_{0}=p_{1}$, we can take the $2^{n}$ dimensional subspace $V_{n}:=2^{-n/{p_{0}}}\cdot\ell^{p_{0}}\{1,\dots,2^{n}\}\hookrightarrow\ell^{q_{0}}(2^{-j/{p_{0}}}\cdot\ell^{p_{0}}\{1,\dots,2^{j}\})_{j=0}^{\infty}$, where $n\geq 1$. The embedding (3.4) restricted to $V_{n}$ is the identity map, hence the Bernstein numbers are all bounded from below by $1$, which does not tend to 0.
In the following we consider $0<p_{0}<p_{1}<\infty$. For each $n\geq 1$ let $S_{n}:=(\{-1,1\}^{n},\mu_{n})$ be the probability space where $\mu_{n}\{a\}=2^{-n}$ for all $a\in\{-1,1\}^{n}$. For $1\leq j\leq n$, let $r_{n,j}:S_{n}\to\{-1,1\}$ be the standard projection on the $j$-th coordinate, i.e. $r_{n,j}(\epsilon_{1},\dots,\epsilon_{n})=\epsilon_{j}$. The standard Khintchine inequality yields $\|\sum_{j=1}^{n}a_{j}r_{n,j}\|_{L^{p}(S_{n},\mu_{n})}\approx_{p}\|a\|_{\ell^{2}\{1,\dots,n\}}$ where the implied constant depends only on $p\in(0,\infty)$ but not on $n$. In particular, there is a $C_{p_{0}p_{1}}>0$ such that for every $n\geq 1$ and $(a_{j})_{j=1}^{n}\subset\mathbb{R}$,
(3.5) $\Big\|\sum_{j=1}^{n}a_{j}r_{n,j}\Big\|_{L^{p_{1}}(S_{n},\mu_{n})}\leq\Big\|\sum_{j=1}^{n}a_{j}r_{n,j}\Big\|_{L^{p_{0}}(S_{n},\mu_{n})}\leq C_{p_{0}p_{1}}\Big\|\sum_{j=1}^{n}a_{j}r_{n,j}\Big\|_{L^{p_{1}}(S_{n},\mu_{n})}.$
Since $(S_{n},\mu_{n})$ and $(\{1,\dots,2^{n}\},2^{-n}\cdot\#)$ are both probability spaces with uniform distributions on their sample sets, we see that they are isomorphic by taking any bijection $\Gamma_{n}:\{1,\dots,2^{n}\}\to S_{n}$ between sets. Thus we have elements $r_{n,j}\circ\Gamma_{n}\in\ell^{p_{1}}\{1,\dots,2^{n}\}$.
Now take $X=\operatorname{span}(r_{n,j}\circ\Gamma_{n})_{j=1}^{n}$, we have $\dim X=n$. By (3.5) we get, for every $(a_{j})_{j=1}^{n}\subset\mathbb{R}$,
$\Big\|\sum_{j=1}^{n}a_{j}\cdot(r_{n,j}\circ\Gamma_{n})\Big\|_{2^{-\frac{n}{p_{0}}}\cdot\ell^{p_{0}}}\leq C_{p_{0}p_{1}}\Big\|\sum_{j=1}^{n}a_{j}\cdot(r_{n,j}\circ\Gamma_{n})\Big\|_{2^{-\frac{n}{p_{1}}}\cdot\ell^{p_{1}}}.$
We conclude that $b_{n}\big(\ell^{q_{0}}(2^{-\frac{j}{p_{0}}}\cdot\ell^{p_{0}}\{1,\dots,2^{j}\})_{j=A}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{-\frac{j}{p_{1}}}\cdot\ell^{p_{1}}\{1,\dots,2^{j}\})_{j=A}^{\infty}\big)\geq C_{p_{0}p_{1}}^{-1}$, which does not go to zero as $n\to\infty$. ∎
Corollary 3.9.
Let $n,A\geq 1$, $0<p_{1}\leq p_{0}\leq\infty$ and $0<q_{0}\leq q_{1}\leq\infty$. Consider the embedding
(3.6) $\ell^{q_{0}}(2^{j(s-n/p_{0})}\cdot\ell^{p_{0}}\{1,\dots,2^{nj}\})_{j=A}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{j(s-n/p_{1})}\cdot\ell^{p_{1}}\{1,\dots,2^{nj}\})_{j=A}^{\infty}.$
  1. (i)
    Suppose $q_{0}<q_{1}$ and $p_{1}<p_{0}=\infty$, then the embedding is finitely strictly singular.
  2. (ii)
    Suppose $q_{0}<q_{1}$, and either $p_{1}\leq p_{0}<\infty$ or $p_{0}=p_{1}=\infty$, then the embedding is strictly singular but not finitely strictly singular.
  3. (iii)
    Suppose $q_{0}=q_{1}$, then the embedding is not strictly singular.
Proof.
The map $J_{s}(x_{j})_{j=A}^{\infty}:=(2^{js}x_{j})_{j=A}^{\infty}$ defines an isomorphism $J_{s}:\ell^{q}(2^{j(s-\frac{n}{p})}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=A}^{\infty}\to\ell^{q}(2^{-j\frac{n}{p}}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=A}^{\infty}$ for every $0<p,q\leq\infty$. Therefore it suffices to consider the case $s=0$.
i: For $j\geq 0$ and $f:\{1,\dots,2^{nj}\}\to\mathbb{R}$ we define $\Gamma_{j}f:[0,1]\to\mathbb{R}$ by $\Gamma_{j}f(x):=f(\lfloor 2^{nj}x\rfloor)$. Clearly $\Gamma_{j}:2^{-jn/p}\cdot\ell^{p}\{1,\dots,2^{nj}\}\to L^{p}[0,1]$ are isometric embeddings for all $0<p\leq\infty$ and $j\geq 0$. Therefore $(\Gamma_{j})_{j=A}^{\infty}:\ell^{q}(2^{-jn/p}\cdot\ell^{p}\{1,\dots,2^{nj}\})_{j=A}^{\infty}\to\ell^{q}(L^{p}[0,1])$ is an isometry for all $0<p,q\leq\infty$ and $A\geq 1$.
Applying Corollary 3.4 we see that $\ell^{q_{0}}(\ell^{\infty}\{1,\dots,2^{nj}\})_{j=A}^{\infty}\to\ell^{q_{1}}(L^{p_{1}}[0,1])$ is finitely strictly singular. Pulling back by the topological embedding $(\Gamma_{j})_{j=A}^{\infty}:\ell^{q_{k}}(2^{-jn/p_{k}}\{1,\dots,2^{nj}\})_{j=A}^{\infty}\to\ell^{q_{k}}(L^{p_{k}}[0,1])$ for $k=0,1$ (see also Lemma 4.1) we conclude that (3.6) is finitely strictly singular.
ii: The non finitely strictly singular part follows from Proposition 3.8. The strictly singular part follows from applying Proposition 3.7 to the second arrow below:
[…]
iii: When $q_{0}=q_{1}$, we consider the infinite dimensional subspace $X=\operatorname{span}(\mathbf{1}_{\{1,\dots,2^{nj}\}}\otimes e_{j}:j\geq A)$. Since $\|\mathbf{1}_{\{1,\dots,2^{nj}\}}\|_{2^{-jn/p}\cdot\ell^{p}}\equiv 1$ for all $j\geq 0$ and $0<p\leq\infty$, the embedding restricted to $X$ is an isometry. Therefore, it has a bounded inverse, which proves the non strictly singularity. ∎
4. Proof of Theorems 1.1 and 1.2
In this section we complete the proof of Theorems 1.1 and 1.2.
Lemma 4.1.
Let $X_{0},X_{1},Y_{0},Y_{1}$ be quasi-Banach spaces, and $\iota_{0}:X_{0}\to Y_{0}$ and $\iota_{1}:X_{1}\to Y_{1}$ be topological embeddings, i.e. $\iota_{0}(X_{0})\subseteq Y_{0}$ and $\iota_{1}(X_{1})\subseteq Y_{1}$ are both closed subspaces. Let $T_{X}:X_{0}\to X_{1}$ and $T_{Y}:Y_{0}\to Y_{1}$ be bounded linear maps such that $\iota_{1}\circ T_{X}=T_{Y}\circ\iota_{0}$.
  1. (i)
    Suppose $T_{Y}$ is (finitely) strictly singular. Then so is $T_{X}$.
  2. (ii)
    Conversely suppose $T_{X}$ is not (finitely) strictly singular. Then so is $T_{Y}$.
Here in the lemma we have the commutative diagram $Y_{0}$$Y_{1}$$X_{0}$$X_{1}$$\scriptstyle{T_{Y}}$$\scriptstyle{T_{X}}$$\scriptstyle{\iota_{0}}$$\scriptstyle{\iota_{1}}$.
Proof.
Since $\iota_{0},\iota_{1}$ are both topological embeddings, there is a $C>0$ such that for $k=0,1$ and for all $x\in X_{k}$, $C^{-1}\|x\|_{X_{k}}\leq\|\iota_{k}x\|_{Y_{k}}\leq C\|x\|_{X_{k}}$. Both results on the finitely strictly singularity follow from the fact that $b_{n}(T_{X})\leq C^{2}b_{n}(T_{Y})$. The strictly singularity follows from the same argument and we leave the details to the reader. ∎
We can decompose the proof of Theorems 1.1 and 1.2 into a few critical cases.
For embeddings between Besov spaces with different levels of smoothness, we have:
Proposition 4.2.
Let $0<p_{0},p_{1},q_{0},q_{1}\leq\infty$ and $s_{0},s_{1}\in\mathbb{R}$ satisfy $s_{0}-s_{1}\geq\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$, and $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain.
  1. (i)
    If $s_{0}-s_{1}>\frac{n}{p_{0}}-\frac{n}{p_{1}}$ or $q_{0}<q_{1}$, then $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$ is finitely strictly singular. In particular $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$ is also finitely strictly singular.
  2. (ii)
    If $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ and $q_{0}=q_{1}$, then $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$ is not strictly singular. In particular $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$ is also not strictly singular.
Remark 4.3.
When $s_{0}-s_{1}>\frac{n}{p_{0}}-\frac{n}{p_{1}}$, the embedding $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\to\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$ is indeed compact, in particular it is finitely strictly singular. The only interesting case on the domain $\Omega$ is the case where $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ and $q_{0}<q_{1}$.
Proof of Proposition 4.2.
Take an $\varepsilon>0$ where $p_{0},q_{0},p_{1},q_{1}>\varepsilon$ and $|s_{0}|,|s_{1}|<\varepsilon^{-1}$. By Proposition 2.13 iv, $\Lambda:\mathscr{B}_{p_{k}q_{k}}^{s_{k}}(\mathbb{R}^{n})\to\ell^{q_{k}}(2^{j(s_{k}-\frac{n}{p_{k}}+\frac{n}{2})}\cdot\ell^{p_{k}}(\mathbb{Z}))_{j=0}^{\infty}$ are isomorphisms for $k=0,1$.
i: By Corollary 3.6 the embedding $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}}+\frac{n}{2})}\cdot\ell^{p_{0}}(\mathbb{Z}))_{j=0}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}}+\frac{n}{2})}\cdot\ell^{p_{1}}(\mathbb{Z}))_{j=0}^{\infty}$ is finitely strictly singular. Therefore $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$ is also finitely strictly singular. Applying Lemma 4.1 with $X_{k}=\mathscr{B}_{p_{k}q_{k}}^{s_{k}}(\Omega)$, $Y_{k}=\mathscr{B}_{p_{k}q_{k}}^{s_{k}}(\mathbb{R}^{n})$ and $\iota_{k}=E|_{X_{k}}$ (see Lemma 2.15) for $k=0,1$, we get that $\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\Omega)\to\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\Omega)$ is also finitely strictly singular.
$i$: When $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ and $q_{0}=q_{1}$, we are going to find an infinite dimensional subspace $X$ such that the embedding restricted to $X$ has bounded inverse.
[…]
$\Big\|\sum_{j=A}^{\infty}a_{j}\psi_{jm}^{g}\Big\|_{\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})}\approx\big\|\big(2^{j(s_{0}-\frac{n}{p_{0}}+\frac{n}{2})}a_{j}\big)_{j=A}^{\infty}\big\|_{\ell^{q_{0}}}=\big\|\big(2^{j(s_{1}-\frac{n}{p_{1}}+\frac{n}{2})}a_{j}\big)_{j=A}^{\infty}\big\|_{\ell^{q_{1}}}\approx\Big\|\sum_{j=A}^{\infty}a_{j}\psi_{jm}^{g}\Big\|_{\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})}.$
By Lemma 2.12 (and a possible change of the implied constants), the same estimate holds if we replace the norm $\|\cdot\|_{\mathscr{B}_{p_{k}q_{k}}^{s_{k}}(\mathbb{R}^{n})}$ by $\|\cdot\|_{\mathscr{B}_{p_{k}q_{k}}^{s_{k}}(\Omega)}$ ($k=0,1$). This completes the proof. ∎
For embeddings between Besov spaces with the same level of smoothness, we have:
Proposition $4.$
Let $s\in\mathbb{R}$, $0<p_{1}\leq p_{0}\leq\infty$ and $0<q_{0}\leq q_{1}\leq\infty$. Let $\Omega\subset\mathbb{R}^{n}$ be a bounded Lipschitz domain. Consider the embedding $\mathscr{B}_{p_{0}q_{0}}^{s}(\Omega)\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s}(\Omega)$.
  1. (i)
    Suppose $q_{0}<q_{1}$ and $p_{1}<p_{0}=\infty$. Then the embedding is finitely strictly singular.
  2. (ii)
    Suppose $q_{0}<q_{1}$, and either $p_{0}<\infty$ or $p_{0}=p_{1}=\infty$. Then the embedding is strictly singular but not finitely strictly singular.
  3. (iii)
    Suppose $q_{0}=q_{1}$. Then the embedding is not strictly singular.
Proof.
Still we take an $\varepsilon>0$ be such that $p_{0},q_{0},p_{1},q_{1}>\varepsilon$ and $|s_{0}|,|s_{1}|<\varepsilon^{-1}$. Let $A\geq 1$ and $R_{j},S_{j}\subset Q_{j}\times G_{j}$ for $j\geq A$ be from Proposition 2.16 where we choose fixed subsets $V\Subset\Omega\Subset U$. Recall that $R_{j}=2^{(j-A)n}$ for $j\geq A$, $S_{j}=2^{(j+A)n}$ for $j\geq 0$, and for $k=0,1$ both arrows in $\ell^{q_{k}}(2^{j(s-\frac{n}{p_{k}}+\frac{n}{2})}\cdot\ell^{p_{k}}(R_{j}))_{j=A}^{\infty}\xrightarrow{\Lambda^{-1}}\mathscr{B}_{p_{k}q_{k}}^{s}(\Omega)\xrightarrow{\Lambda\circ E}\ell^{q_{k}}(2^{j(s_{k}-\frac{n}{p_{k}}+\frac{n}{2})}\cdot\ell^{p_{k}}(S_{j}))_{j=0}^{\infty}$ are bounded.
Denote $J:\mathscr{B}_{p_{0}q_{0}}^{s}(\Omega)\hookrightarrow\mathscr{B}_{p_{1}q_{1}}^{s}(\Omega)$, $I_{R}:\ell^{q_{0}}(2^{j(s-\frac{n}{p_{0}}+\frac{n}{2})}\cdot\ell^{p_{0}}(R_{j}))_{j=A}^{\infty}\to\ell^{q_{1}}(2^{j(s-\frac{n}{p_{1}}+\frac{n}{2})}\cdot\ell^{p_{1}}(R_{j}))_{j=A}^{\infty}$ and $I_{S}:\ell^{q_{0}}(2^{j(s-\frac{n}{p_{0}}+\frac{n}{2})}\cdot\ell^{p_{0}}(S_{j}))_{j=0}^{\infty}\to\ell^{q_{1}}(2^{j(s-\frac{n}{p_{1}}+\frac{n}{2})}\cdot\ell^{p_{1}}(S_{j}))_{j=0}^{\infty}$ for convenience. Recall from Corollary 3.9 we see that for i, ii and iii respectively:
  • Suppose $q_{0}<q_{1}$ and $p_{1}<p_{0}=\infty$. Then $I_{S}$ is finitely strictly singular.
  • Suppose $q_{0}<q_{1}$, and either $p_{0}<\infty$ or $p_{0}=p_{1}=\infty$. Then $I_{S}$ is strictly singular and $I_{R}$ is not finitely strictly singular.
  • Suppose $q_{0}=q_{1}$. Then $I_{R}$ is not strictly singular.
The results then follow from applying Lemma 4.1 with $(T_{X},T_{Y},\iota)=(I_{R},J,\Lambda^{-1})$ and $(J,I_{S},\Lambda\circ E)$. ∎
Proof of Theorem 1.1.
i and ii follow from Corollary 2.18 ii.
iii: By Corollary 2.18 $i$ the map is noncompact. For finitely strictly singularity, when $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ the result follows from Proposition 4.2 i; when $s_{0}-s_{1}=0$ the result follows from Proposition 4.4 i.
iv: The result follows from Proposition 4.4 ii.
v: When $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}$ the result follows from Proposition 4.2 $i$; when $s_{0}-s_{1}=0$ the result follows from Proposition 4.4 iii. ∎
Proof of Theorem 1.2.
i follows from Corollary 2.18 i.
$i$: By Corollary 2.18 i the map is noncompact. The finitely strictly singularity follows from Proposition 4.2 i.
iii: When $p_{0}>p_{1}$, i.e. the case $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$ and $q_{0}=q_{1}$, the result follows from Proposition 4.2 ii. The only remaining case left from the previous discussion is the case $p_{0}=p_{1}$. Recall that either $s_{0}-s_{1}>0$, or $s_{0}-s_{1}=0$ but $q_{0}\leq q_{1}$.
Indeed, choose an $\varepsilon>0$ such that $p_{0}(=p_{1}),q_{0},q_{1}>\varepsilon$ and $|s_{0}|,|s_{1}|<\varepsilon^{-1}$. Let $\psi_{jm}^{g}$ be the wavelet elements in Proposition 2.13. We can consider the infinite dimensional subspace $X=\operatorname{span}\{\psi_{0m}^{0}:m\in\mathbb{Z}^{n}\}$. By Proposition 2.13 iv, for every $(a_{m})_{m\in\mathbb{Z}^{n}}\subset\mathbb{R}$,
$\Big\|\sum_{m\in\mathbb{Z}^{n}}a_{m}\psi_{0m}^{0}\Big\|_{\mathscr{B}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})}\approx\|(a_{m}\big)_{m\in\mathbb{Z}^{n}}\|_{\ell^{p_{0}}}=\|(a_{m}\big)_{m\in\mathbb{Z}^{n}}\|_{\ell^{p_{1}}}\approx\Big\|\sum_{m\in\mathbb{Z}^{n}}a_{m}\psi_{0m}^{0}\Big\|_{\mathscr{B}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})}.$
Thus the embedding restricted to $X$ has bounded inverse, completing the whole proof. ∎
5. On Homogeneous Besov Embeddings
As a bonus result to Theorem 1.2, we state and prove the analogous classification result on homogeneous Besov spaces by reviewing the methods used in the previous discussion.
Recall that the homogeneous Besov space $\dot{\mathscr{B}}_{pq}^{s}(\mathbb{R}^{n})$ is defined to be the element in the quotient space $\dot{\mathscr{S}}^{\prime}(\mathbb{R}^{n})=\mathscr{S}^{\prime}(\mathbb{R}^{n})/\{\text{polynomials}\}$ for which
$\|f\|_{\dot{\mathscr{B}}_{pq}^{s}(\mathbb{R}^{n})}=\|f\|_{\dot{\mathscr{B}}_{pq}^{s}(\mathbb{R}^{n};\varphi)}:=\Big(\sum_{j\in\mathbb{Z}}\|2^{js}\varphi_{j}\ast f\|_{L^{p}(\mathbb{R}^{n})}^{q}\Big)^{1/q}<\infty.$
Here $(\varphi_{j})_{j\in\mathbb{Z}}\subset\dot{\mathscr{S}}(\mathbb{R}^{n})$ is a fixed family of Schwartz functions with all moment vanishings, such that $\operatorname{supp}\hat{\varphi}_{0}\subset\{\frac{1}{2}<|\xi|<2\}$, $\hat{\varphi}_{j}(\xi)=\hat{\varphi}_{0}(2^{-j}\xi)$ for all $j\in\mathbb{Z}$, and $\sum_{j\in\mathbb{Z}}\hat{\varphi}_{j}(\xi)=1$ for all $\xi\neq 0$. See e.g. [Tri10, Chapter 5] for details.
Theorem 5.1.
Let $p_{0},p_{1},q_{0},q_{1}\in(0,\infty]$, $s_{0},s_{1}\in\mathbb{R}$. Consider the embedding $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
  1. (i)
    $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\not\subset\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$, i.e. there is no embedding, if and only if one of the following occurs:
    • $s_{0}-s_{1}\neq\frac{n}{p_{0}}-\frac{n}{p_{1}}$;
    • $p_{0}>p_{1}$ (i.e. $\frac{n}{p_{0}}-\frac{n}{p_{1}}<0$) or $q_{0}>q_{1}$.
  2. (ii)
    The embedding is non-compact but finitely strictly singular iff $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}>0$ and $q_{0}<q_{1}$.
  3. (iii)
    The embedding is not strictly singular if and only if one of the following occurs:
    • $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}=0$ and $q_{0}\leq q_{1}$;
    • $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}\geq 0$ and $q_{0}=q_{1}$.
[…]
Proof.
Firstly the wavelet decomposition holds for homogeneous function spaces, see e.g. [FJW91, Theorem 6.16]: there is a collection $\{\dot{\psi}_{jv}:j\in\mathbb{Z},v\in\mathbb{N}_{0}\}\subset\dot{\mathscr{S}}(\mathbb{R}^{n})$ which forms an orthonormal basis to $L^{2}(\mathbb{R}^{n})$, such that for every $f\in\dot{\mathscr{S}}^{\prime}(\mathbb{R}^{n})$ we have decomposition $f=\sum_{j\in\mathbb{Z}}\sum_{v=0}^{\infty}\dot{(}f,\dot{\psi}_{jv})\dot{\psi}_{jv}$ with the summation converges in $\dot{\mathscr{S}}^{\prime}(\mathbb{R}^{n})$. Moreover, $\Lambda(f):=\{(f,\dot{\psi}_{jv}):v\in\mathbb{N}_{0}\}_{j\in\mathbb{Z}}$ defines a map which gives the isomorphisms
$\Lambda:\dot{\mathscr{B}}_{pq}^{s}(\mathbb{R}^{n})\xrightarrow{\simeq}\ell^{q}(2^{j(s-\frac{n}{p})}\cdot\ell^{p})_{j=-\infty}^{\infty},\qquad\text{for all }0<p,q\leq\infty\text{ and }s\in\mathbb{R}.$
Therefore it is equivalent to consider the embedding $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}})_{j=-\infty}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}})_{j=-\infty}^{\infty}$.
i: In this assumption $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}})_{j=-\infty}^{\infty}\not\subset\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}})_{j=-\infty}^{\infty}$, thus $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\not\subset\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
To be more precise, when $s_{0}-s_{1}\neq\frac{n}{p_{0}}-\frac{n}{p_{1}}$ the result follows from that $2^{j(s_{0}-\frac{n}{p_{1}})}\big/2^{j(s_{0}-\frac{n}{p_{0}})}\to\infty$ for either $j\to+\infty$ or $j\to$-$\infty$.
When $s_{0}-s_{1}=\frac{n}{p_{0}}-\frac{n}{p_{1}}=:s$, the map $J_{s}(x_{j})_{j\in\mathbb{Z}}:=(2^{js}x_{j})_{j\in\mathbb{Z}}$ defines isomorphisms that for $i=0,1$, $J_{s}:\ell^{q_{i}}(2^{js}\cdot\ell^{p_{i}})_{j\in\mathbb{Z}}\to\ell^{q_{i}}(\mathbb{Z};\ell^{p_{i}})$ . Since $\ell^{q_{0}}(\ell^{p_{0}})\not\subset\ell^{q_{1}}(\ell^{p_{1}})$ holds when $q_{0}>q_{1}$ or $p_{0}>p_{1}$, pulling back from $J_{s}$ we get the result.
ii: In this assumption, applying Proposition 3.5 the map $\ell^{q_{0}}(\mathbb{Z};\ell^{p_{0}})\hookrightarrow\ell^{q_{1}}(\mathbb{Z};\ell^{p_{1}})$ is finitely strictly singular. Clearly it is non-compact. Pulling back from $J_{s}$, we get the non-compactness and finitely strictly singularity of $\ell^{q_{0}}(2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}})_{j=-\infty}^{\infty}\hookrightarrow\ell^{q_{1}}(2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}})_{j=-\infty}^{\infty}$, so is $\dot{\mathscr{B}}_{p_{0}q_{0}}^{s_{0}}(\mathbb{R}^{n})\hookrightarrow\dot{\mathscr{B}}_{p_{1}q_{1}}^{s_{1}}(\mathbb{R}^{n})$.
iii: When $q_{0}=q_{1}$, the result follows from Lemma 3.1; when $p_{0}=p_{1}$ the result from the fact that for every $j\in\mathbb{Z}$, the spaces $2^{j(s_{0}-\frac{n}{p_{0}})}\cdot\ell^{p_{0}}=2^{j(s_{1}-\frac{n}{p_{1}})}\cdot\ell^{p_{1}}$ and have the same norms. ∎
6. Further Remarks
We conclude our paper with a few remarks and open questions.
In Theorems 1.1 and 1.2, we observe the impact of differences between function spaces on bounded Lipschitz domains $\Omega$ and on $\mathbb{R}^{n}$. A natural question that arises is what results can be obtained for embeddings on unbounded open sets (say unbounded uniform domains).
We believe that the upper bounds in Proposition 3.5 could be sharpened, and that it should be possible to derive corresponding lower bounds. This could lead to improved estimates on the rate of decay for Bernstein numbers between Besov spaces when the embedding is non-compact but finitely strictly singular. Note that although we do not have a canonical definition of Besov norms, the decay rates of their Bernstein numbers are determined up to a constant multiple which still make sense. The case where $\Omega=\mathbb{R}^{n}$ is particularly intriguing, which is equivalent to computing the decay rate of $b_{n}\big(\ell^{q_{0}}(\ell^{p_{0}})\to\ell^{q_{1}}(\ell^{p_{1}})\big)$.
Given the obtained results for Besov spaces, it seems natural to ask about the inner structure for non-compact embeddings between Triebel-Lizorkin spaces. This could be quite interesting, given their different behaviors of the corresponding sequence spaces from the wavelet decomposition, which makes the study of Triebel-Lizorkin spaces more challenging. Also, it is worth noting that the obtained results for Besov spaces could be used for non-compact Sobolev embeddings which will generalize and extend results from [LM23] and [BG89].
Acknowledgment.
The author thanks Luis Rodríguez-Piazza for some helpful discussion.